OpenStudy (anonymous):

Does ∑ 1 to ∞ (n^2 ÷(e^n /3) ) converge or diverge?

OpenStudy (anonymous):

You need to apply a test for converge, which one that you know of works here?

OpenStudy (anonymous):

I'm thinking Limit Comparision Test but I'm not sure which series would work to compare it with

OpenStudy (anonymous):

I tried e^n/3 but that doesn't work

OpenStudy (anonymous):

hmm, that's not what I had in mind. It might work if you choose the other series properly, but I'd try a different test.

OpenStudy (anonymous):

Which test do you think would work?

OpenStudy (anonymous):

the root test.

OpenStudy (anonymous):

I know the root test but this problem was in a section in the book before the root test (by then, there were only nth term test, DCT, LCT, Integral Test) so it should be solvable by one of these tests

OpenStudy (anonymous):

Ok, the root test would be a lot easier, but the integral test might work.

OpenStudy (anonymous):

I haven't thought about solving it using the intergral test really, so I'll try that I guess and see where I get

OpenStudy (anonymous):

You need to integrate by parts twice, so have fun.

OpenStudy (anonymous):

Yep, that works

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty} \frac{k^2}{\frac{e^k}{3}}=3\sum_{k=1}^{\infty}\frac{k^2}{e^k}\] There exists m (natural number) such that m < k implies k^4 < e^k. Then\[3\sum_{k=1}^{\infty}\frac{k^2}{e^k}=3(\sum_{k=1}^{m}\frac{k^2}{e^k}+\sum_{k=m+1}^{\infty}\frac{k^2}{e^k})\le 3(\sum_{k=1}^{m}\frac{k^2}{e^k}+\sum_{k=1}^{m}\frac{k^2}{k^4})\] In the expression, the first term is finite (a definite sum), the second converges. Convergence by comparison test.

OpenStudy (anonymous):

Limits of summation on the last term should be k=m+1 to infinity. Sorry for the shoddy attention to detail....